### v1

# TRIM: Token Reduction and Inference Modeling for Cost-Effective Language Generation

Abstract
The inference cost of Large Language Models (LLMs) is a significant challenge due to their computational demands, specially on tasks requiring long outputs. However, natural language often contains redundancy, which presents an opportunity for optimization. We have observed that LLMs can generate distilled language—concise outputs that retain essential meaning, when prompted appropriately. We propose a framework for saving computational cost, in which a shorter distilled output from the LLM is reconstructed into a full narrative by a smaller model with lower inference costs. Our experiments show promising results, particularly in general knowledge domains with 20.58% saved tokens on average with tiny decrease in evaluation metrics, hinting that this approach can effectively balance efficiency and accuracy in language processing tasks.
TRIM: Token Reduction and Inference Modeling for Cost-Effective Language Generation
Alfredo Garrachón Ruiz  and Tomás de la Rosa  and Daniel Borrajo AI Research, JPMorganChase
1 Introduction
Large language models (LLMs) have shown remarkable capabilities across a wide range of tasks, from natural language understanding to creative content generation. However, the computational cost of inference and the associated energy consumption present significant challenges. As the demand for AI applications continues to grow, these costs are expected to escalate, raising concerns about sustainability and accessibility Wu et al. (2022). Addressing these issues is crucial to ensure that future advancements in AI remain feasible and widely deployable. Hence, we propose a new technique to save tokens at inference time skipping less informative and semantic-irrelevant function words. Later on, we propose to regenerate them based on the context with a cost-effective alternative, leveraging on the feature that natural language is considerable redundant.
As part of our feasibility analysis, we empirically verified that this task is achievable by state-of-the-art LLMs, even though they are not specifically designed for such tasks, and the altered output may not have been encountered in any part of the training corpus. Given that LLMs can generate this distilled output, the omitted words can be reinserted into the output text using a language model with a low inference cost. Studies in cognitive linguistics suggest that human readers naturally infer missing function words based on context Ferreira and Patson (2007). Similarly, transformer-based models use masked language modeling, where parts of a sentence are intentionally masked, and the model learns to predict the masked tokens based on context Devlin et al. (2019). Our hypothesis is that small language models can mirror this human reading abilities of filling the gaps for low-information inferable terms.
In this work, the main contribution is the proposal of a novel technique for reducing the inference costs of LLMs through the suppression and later-on reconstruction of function words. This approach is orthogonal to other optimization techniques, and could be applicable as LLMs continue to grow in size and capabilities. We also propose an algorithm to check and define the applicability of this technique in different domains, selecting the most proper function words set, and analyzing the lose in performance as the percentage of saved tokens increases. Additionally, we provide an experimental evaluation in the context of general knowledge question-answering. In this setting, the output length is more relevant than the input query, making the applicability of our technique more evident. The results show that even suppressing around 18% of the full text, models that are hundreds of times smaller can reconstruct distilled text without losing significant text quality.
2 Related Work
In this section we review several approaches to optimize LLM inference which are related to our work.
2.1 Prompt Compression
One alternative to decrease the computational cost is setting the focus on reducing the input by prompt compression. This could be a token-level removal of less relevant words Jiang et al. (2023); Pan et al. (2024); Fu et al. (2024), or more elaborated approaches such as context-aware sentence encoding Liskavets et al. (2024). This task operates on the input rather than on the output. Therefore, our approach would be complementary to prompt compression. Moverover, prompt compression does not require a later reconstruction of original input text, as it pertains to the encoding step and the result of the compression is handled internally with vectorized representations.
2.2 Inference Optimizations
For the generative-LLM inference, two phases are recognizable. In the first one, all the input tokens from the prompt are computed in parallel. In the second one, tokens are generated in successive passes using previously generated tokens, and the key-value cache collected in the first phase Miao et al. (2023a). Given the memory-intense requirement, the token generative phase might be allocated en specialized hardware Patel et al. (2024). On the other hand, smaller LLM models can be used to generate multiple output sequences in parallel that then are merged and verified Miao et al. (2023b). Another way to save costs at inference time consists of building a sequence of increasing-size LLMs, where a (larger) model at step $N$ is only queried when the answer of the model at step $N-1$ is not deem reliable Chen et al. (2023).
2.2.1 Text Reconstruction
One of the core learning ideas in Natural Language Processing (NLP) is having a vectorized representation that allows models to reconstruct back the original text Oshri and Khandwala (2015). Nevertheless, NLP tasks, such as masking Devlin et al. (2019) or in a more general scope, infilling Donahue et al. (2020), rely on markers indicating where to substitute the text. In our approach, the input text at inference time would not have indication of what positions contained an omitted word. Text with omitted words can also be used as negative examples, to improve the performance of automatic translators Yang et al. (2019), but to our knowledge omitted/original texts have not been used to train for inference efficiencies.
3 Exploratory Research
The feasibility of our approach relies on the LLM being able to perform the task of generating an answer omitting a set of terms that do not contribute to the semantic meaning. This is a task for which the LLM was not originally trained. To verify this assumption, we performed an exploratory analysis where we used the Gpt-4o model OpenAI (2024), and developed a prompt to carry out this task indicating also that omitted terms should be easily inferrable in a later stage. The prompt includes a pre-defined set of terms composed by determiners and conjunctions with a total of 23 terms such as: (the, a, an, and, but, so, that). See supplementary material (Section A.1) for details. We asked the LLM to generate the answer to 500 knowledge questions (See dataset in Section 6) with and without the previously defined prompt. We computed the average number of terms from the set present in both types of answers (Table 1). We observed how the LLM is able to complete the task effectively, generating the answer omitting nearly all terms. Additionally, the average is not zero, indicating that in some scenarios the LLM preserves some of these terms pro-actively deciding that they are not easily reconstructible a posteriori in that scenario.
Text Type Avg. count from word set Std. Deviation
Original 22.50 5.77
Distilled 0.32 0.62
Table 1: Average count of the terms present in the text.
4 Semantic Relevance Distribution Of Words In The English Language
The importance of some terms when measuring the semantic meaning of texts is a well-documented phenomenon, particularly in English. A study by Zipf (1949) highlighted that "speakers tend to use more common words more frequently to reduce the cognitive and physical effort involved in communicatingZipf (1949). Certain words, such as articles and determiners, often contribute little to the overall semantic meaning of a sentence. These words are frequently classified as function words, which serve a grammatical purpose rather than adding significant content or meaning to a sentence Jurafsky and Martin (2000); Manning and Schütze (1999).
Even though function words provide little semantic meaning, not all of them can be effectively reconstructed on the basis of context. Some of them, such as pronouns or quantifiers, cannot be always inferred based on context. For example, in the sentence "I went to the marathon in the city center", if we remove the function words, the sentence will be "went marathon city center". The terms "the", "to", and "in" are easily recovered based on the context, but the first pronoun could be any of the personal pronouns.
The categorization and role of function words are clear from the English grammar and linguistic theory. However, few NLP studies have analyzed them regarding how they affect representation learning Kim et al. (2019). Besides, our interest in this paper lies on how easy it is for a language model to infer the function words, so then we can pre-select the best set of words to include in the instruction prompt. For this purpose, we propose a new algorithm based on the use of language models trained with the masking technique that allows to obtain a sorted list of the most easily inferable words based on the context. The algorithm focuses on calculating the difference in probability between the actual word and the next most probable word in a fixed position based on the context. This difference represents the prevalence of the actual word over the distribution of the rest of possible words. This means that, the bigger the difference, the less doubt a language model has about choosing that term over the rest, and, therefore, it is easily inferable based on the context.
For a given text $c$ (e.g. a paragraph from a corpus) and a word $w$, let $c^{\prime}$ be the text $c$ with the term $t$ masked. We define the probability difference as:
$\Delta P=P(w\mid c^{\prime})-P(w_{\rm{alt}}\mid c^{\prime})$
where $P(w_{\rm{alt}}\mid c)$ is the probability of the most probable alternative term ($w_{\rm{alt}}\neq w$) given the same context, obtained using the LM. The sorted list of function words based on this criteria is computed following Algorithm 1. After computing $\Delta P$ for all function words in each corpus fragment, it computes the average for each word, and returns the list, sorted by this average.
Algorithm 1 Terms sorted by ease of inference
1:Input: Word set $T=\{w_{1},w_{2},\dots,w_{n}\}$,
2:   text corpus $C=\{c_{1},c_{2},\dots,c_{m}\}$,
3:   language model $LM$
4:Output: Sorted T by probability difference $\Delta P_{j}$
5:for each $c_{i}\in C$ do
6:     for each $t_{j}\in T$ do
7:         if $t_{j}\in c_{i}$ then
8:              Compute $\Delta P_{i,j}$ with $LM$
9:         end if
10:     end for
11:end for
12:for each $t_{j}\in T$ do
13:     $\Delta P_{j}=$ average$(\Delta P_{i,j})$
14:end for
15:$T^{\prime}$ = Sort $T$ by $\Delta P_{j}$ in descending order
16:return T’
5 Cost-Saving Pipeline
The exploratory research showed the feasibility of the initial idea. Now, we cover the development and specification of the approach. We propose a pipeline for wrapping the use of generative LLMs (both closed and open) to save costs during output generation through the philosophy of omitting the generation of inferable function words, and reconstructing the full text using a much smaller and optimized language model later on. This pipeline can be seen in Figure 1. It can be defined by the following elements:
Inferable terms: Let $S$ be a set of words deemed as easily inferable based of the context (i.e. stop words, functions words…). This set can be experimentally collected from a corpus using Algorithm 1 proposed earlier.
Generation model: Let $M_{G}$ be the generative LLM model capable of generating a response to a query, omitting terms from $S$. The response is called distilled answer $A_{D}$.
Reconstructor model: Let $M_{R}$ be an encoder-decoder model fine-tuned to reconstruct truncated output into a full text with regards of the specific set of terms $S$.
Instruction Prompt: Let $\mathcal{P}$ be the prompt designed to instruct the LLM in the task of omitting words given $S$.
Being $Q$ the query from the user and $A$ the expected answer returned to the user in normal conditions, the pipeline process could be exemplified as:
$A_{D}=M_{G}(Q,P(S))$
$A=M_{R}(A_{D})$
Figure 1: Diagram of the proposed methodology against a vanilla technique.
6 Datasets & Models
For the experimentation, we used two datasets. First, the Wikipedia dump, which is composed of a large amount of well-written and grammatically correct data on the domain of knowledge Foundation . Different subsets from it will be used for analyzing the ease of inferring functions words, and also for training and testing reconstruction models.
Second, we used a synthetic knowledge question-answering dataset we have created for this project. We required an evaluation that aligns with the proposed methodology where the reconstruction model reconstructs distilled text generated by an LLM. For this purpose, we generated 9,000 knowledge questions on different topics. For the generation, we took 30 general knowledge topics, 20 concepts on each of them and 5 keywords associated to each concept, creating various questions from templates regarding explanation, summarization and concept comparison. See supplementary material (Section A.2) for specifications, and an example of these questions.
Later, for each question, GPT-4o was requested to generate both the answer and the distilled version of the answer. To achieve this, we developed the instruction prompt $\mathcal{P}$ that compacts this request with the desired term set $S$, following the few-shot philosophy. Ideally, the generation of the original and distilled answer should be requested in different sessions, so the distilled one would simulate the real scenario of our pipeline, and the original would be the ground truth. Otherwise, the LLM would have the context of the original to generate the distilled version. However, after carrying some experiments we realized that doing it in different sessions produce additional variability (e.g., synonyms, grammatical alternative and paraphrasing) that prevent us to use them in the evaluation (see Supplementary Material A.3). Under the assumption that the LLM can generate distilled text directly as shown in Section 3, we keep the single session generation to focus the evaluation on the ability of reconstructing the distilled narrative. Thus, this second dataset called from now on DKA dataset - Distilled Knowledge Answers, is composed of distilled answers with their original answers that only differ on omissions from the input $S$.
On the side of the reconstructor model $M_{R}$, we consider (1) BART, Facebook AI’s model recognized as one of the early transformer-based models with an auto-regressive decoder for text generation Lewis et al. (2019); (2) T5 from Google, which has demonstrated strong performance in summarization and translation tasks Raffel et al. (2023) and finally, (3) Qwen2.5 model from Alibaba, specifically the 0.5B and 1.5B variants, as it was positioned among the top instruct model at its size at the time of comparison, according to the Open LLM Leaderboard Yang et al. (2024) Fourrier et al. (2024). All these models are at least two orders of magnitude smaller than state-of-the-art LLMs, which makes them good candidates to achieve the reconstruction with lower computational costs. Table 2 depicts model size in billion parameters for these models.
BART T5 Qwen2.5-0.5 Qwen2.5-1.5
0.139 0.223 0.5 1.5
Table 2: Size in billions of the selected models.
7 Evaluation
The evaluation is divided into two experiments. The first one applies Algorithm 1 for the identification of a good set of inferable terms $S$ without notable downgrade in the metrics. In the second experiment, once the set $S$ has been fixed, several language models will be trained and evaluated in the task of the reconstructor model $M_{R}$. We also analyzed if the results validate the applicability of the complete pipeline without noticing relevant deterioration in the output.
For the evaluation, we thought it was necessary to find a specific metric that would evaluate the recovery of the terms from the inferable set. The common NLP metrics focus on comparing two pieces of text that can come in different sizes (i.e., word, sentence, paragraph…), but they do not offer the versatility of focusing only on a specific set of terms according to their exact position, covering the cases of omission, substitution, matching or insertion. The Needleman-Wunsch algorithm, used in bio-informatics to align protein or nucleotide sequences, offers these desired features Needleman and Wunsch (1970), so we adapted it to align the paragraphs of the original and distilled text to effectively maintain the alignment over scenarios where there are missing or extra terms. After applying the algorithm, we can compare each of the positions in the alignment and cover the cases mentioned above, obtaining precision, recall and F1 values with regard to the inferable set. From now on, these metrics will be referred as $\Theta$ metrics.
The rest of evaluation metrics employed compare the whole texts and encompass three primary aspects: grammatical similarity, semantic similarity, and coherence. Grammatical similarity between two texts is assessed using the metrics SacreBLEU, METEOR, and ROUGE, which primarily perform n-gram-based comparisons Post (2018); Banerjee and Lavie (2005); Lin (2004). For semantic similarity, we use cosine similarity, which measures the similarity between two texts using BERT embeddings Devlin et al. (2019); Salton et al. (1975). To evaluate textual coherence, we utilize the perplexity metric, which quantifies coherence and fluency based on how easily a language model (in our case, GPT-2 Radford et al. (2019)) can predict the next tokens, where lower values indicates higher fluency and coherence Manning and Schütze (1999). These experiments were performed on an AWS "g4dn.xlarge" machine with an Nvidia "T4" GPU with 16GB of graphics memory.
7.1 Inferable terms set
In this study, we analyze the degradation of the metrics as the set of function words to be distilled and reconstructed increases. This will allow us to empirically decide a good choice of set $S$ for the second experiment. We applied Algorithm  1 using a random sample of 100,000 paragraphs from the Wikipedia dump and a extended list of 127 function words of the English language Jurafsky and Martin (2000).
To determine the boundary for the subset of words that will be reconstructed a posteriori $S$, we followed an iterative procedure in which the T5 model is fine-tuned for the text reconstruction task using the top-N inferable terms from the sorted list, increasing the subset in 5 words, from size 5 up to 60 (12 levels). For this training we used a new random sample from the Wikipedia dump. Each training set consists of 100,000 (original, distilled) paragraph pairs where the distilled text is built by removing the words of the corresponding level. Additional 5,000 pair for testing were built with the same schema for each of the 12 levels. The performance is then evaluated at each of these distillation levels, analyzing the degradation of the results as the set size increases.
The results per level can be seen in Figure 2, representing the values of the metrics against the percentage of saved tokens per level, and extrapolating trend lines. We can observe that from level 5 onward (around 18% of saved tokens), the percentage of tokens saved increases to a lesser extent while the metrics continue to degrade. Therefore, we set our boundary at level 5, and collect these 25 function words (set $S$) for the final evaluation.
Figure 2: Evaluation at different levels of saved tokens.
7.2 Reconstruction model
Once we have segmented the set of inferable terms for this experimentation, we can focus on the development of the reconstruction model $M_{R}$ in charge of inferring the terms omitted from the generated text. For this purpose, we have taken a new random sample of 100,000 paragraphs from the Wikipedia dump, and we have removed the inferable terms of set $S$, thus forming the associated training corpus. The 4 language models selected have been fine-tuned in the text reconstruction task. The training information can be found in Table 4 of the Supplementary Material.
We tested each model on two test sets. First, on a new random sample of the Wikipedia dump to validate the training, and second, on the DKA dataset (Section 6). We evaluated their performance in a scenario closer to the real one where these models are used after the generation of a text omitting the inferable terms by an LLM. Both test datasets are composed of 9,000 paragraphs each. The results can be seen in Table 5 and 3, showing also the metrics of the distilled text compared to the original one.
Regarding the $\Theta$ metrics, we observe good results overall. The smaller language models Bart and T5 excel in precision, and the instruction-bigger language model Qwen2.5 in its 0.5B and 1.5B versions excel in recall. This means the smaller models tend to be more conservative but more precise, while on the other hand the bigger models reconstruct more terms from the original structure but with higher variability in their correctness. Nevertheless, the metrics are very equated, and do not reflect a high difference between them.
Regarding the rest of metrics, we can observe that the grammatical metrics (BLEU, METEOR & ROUGE) have good values across all models, the semantic metric is perfect with 3 decimals approximation, and the perplexity suffers only a small increase over the original one. All of these results are obtained with a significant percentage of saved tokens of 18.85% in the sample of Wikipedia and 20.58% for the answers of the LLM.
We can observe from the metrics for the distilled text how the grammatical metrics are substantially worse, and the perplexity much higher while the semantics metrics remain almost perfect. This shows that the elimination of the inferable sets breaks the grammatical structure and coherence but maintains almost in its fullness the original semantic information.
Focusing on the differences between test sets (DKA dataset and Wikipedia sample), we observe that the results on the DKA dataset are overall better, showing that the LLM is able to perform the task of omitting inferable terms very effectively, and lead to an easy reconstruction. Another interesting insight can be seen in the $\Theta$ metrics of the distilled text (i.e., matching distilled text with the ground thruth). In the Wikipedia sample evaluation, the $\Theta$-recall is 0 as we remove them directly from the text while in the DKA dataset evaluation the $\Theta$ recall is $8.6\%$, showing that the LLM does not erase all of the terms from the list, but only those it considers are easily inferable from the context. This result is also consistent with the exploratory analysis result shown in Table 1.
We do not see a significant difference among models, and any of them can be used effectively for this task. More closely, we observe how the T5 model obtains better results overall than Qwen2.5 in its 0.5B version, and are very comparable with its 1.5B. It beats the 1.5B version in some metrics even being nearly seven times smaller. For its smaller size, associated cost and overal good performance, it seems to be the sweet spot for this reconstruction model.
Bart T5 Qwen2.5-0.5 Qwen2.5-1.5 Distilled Text
$\Theta$ Precision (%) 91.50(6.98) 93.35(6.10) 88.28(8.27) 90.94(7.21) 96.85(14.92)
$\Theta$ Recall (%) 86.18(8.67) 87.14(8.60) 88.93(7.67) 92.04(6.53) 8.61(10.72)
$\Theta$ F1 (%) 88.55(6.92) 89.97(6.67) 88.44(7.13) 91.36(6.11) 14.25(15.76)
SacreBLEU (%) 85.09(8.10) 87.95(7.57) 86.76(7.80) 88.96(7.04) 49.16(11.59)
METEOR (%) 93.69(4.33) 94.65(4.00) 94.66(4.24) 95.59(3.62) 75.42(6.85)
ROUGE-1 (%) 95.99(2.51) 96.66(2.31) 96.00(2.81) 96.64(2.44) 85.86(4.10)
ROUGE-L(%) 94.95(3.18) 95.85(2.88) 94.95(3.47) 95.94(2.94) 85.86(4.10)
Cosine Similarity (%) 99.99(0.01) 100.00(00.01) 99.99(00.02) 99.99(00.02) 99.92(00.04)
Perplexity 19.35(6.83) 17.04(6.10) 18.42(6.68) 17.59(6.23) 99.77(61.36)
Perplexity Original 14.58(4.83) 14.58(4.83) 14.58(4.83) 14.58(4.83) 14.58(4.83)
Saved Tokens (%) 20.58(5.60) 20.58(5.60) 20.58(5.60) 20.58(5.60)
Training Time (h) 1.5 3 7.5 18
Parameters (B) 0.139 0.223 0.5 1.5
Table 3: Average (standard deviation) metrics results over the Distilled Knowledge Answers dataset (DKA).
8 Discussion
8.1 Cost Saving
Estimating the cost of the use of LLM API services at token level is very difficult GPT . GPT-x costs depends on many factors, including type of task, prompt quality, customization, API call volume, and computational resources. Nevertheless, the model size is the most relevant feature when considering model inference cost. Therefore, we want to make an exercise to illustrate how we should think about the potential cost savings of our pipeline. Our proposal should consider the gain/losses of processing text, which for cost purposes is measured by its size. The relevant text fragments are:
Extra inputs ($I$) : the cost of extending the prompt to specify the task of omitting function words. Specifically for our pipeline, the difference between size($Q$) and size($Q,\mathcal{P}(S)$).
Gained output ($G$): the gains from saved tokens compared to the vanilla generation. This is the difference between size($A$) and size($A_{D}$).
Reconstruction inputs ($I_{R}$): the cost of querying for the reconstruction task. This is equivalent to size($A_{D}$).
Reconstruction output ($O_{R}$): the cost of generating the final output. This is equivalent to size($A$).
On the other hand, we express the input/output costs per size unit using the same subscripts as models. So, let $C^{i}_{D},C^{o}_{D}$ be the input and output cost for the generation model, and $C^{i}_{R},C^{o}_{R}$ the input and output cost for the reconstruction model. The requirement framework for which our pipeline application would make sense is expressed by the formula:
$C^{o}_{D}G\geq C^{i}_{D}I+C^{i}_{R}I_{R}+C^{o}_{R}O_{R}$
Assuming the costs of processing text for the reconstruction models, are two orders of magnitude cheaper, as seen in the evaluation, the relevant term in the right hand side is $C^{i}_{D}I$. Using available prices for the GPT-4o API service in its "gtp-4o-2024-08-06" version (i.e., \$2.5 per million input tokens, \$10 per million output tokens) as a proxy of cost, we see that the ratio $C^{i}_{D}/C^{o}_{D}$ is $\sim 0.25$. Therefore, any unrestricted domain knowledge answer will compensate processing $I$, which in our experiments corresponds to 97 extra words (See prompt 2 in Supplementary Material A.1).
8.2 Other language opportunities
Even though this study only focused on English, the technique is language independent. Several studies indicate that most languages also follow the Zipf distribution Piantadosi (2014), and that frequency ranks have a high correlation among languages Calude and Pagel (2011). However, these analysis were interested in word semantics rather than in function words. Moreover, our final inferable set $S$ only has four words in common with the first 100 words from the Swadesh vocabulary Swadesh (1952), a benchmark word list for language comparison. So, the particular benefits depending on the language is an open question, and we leave it as future work. We expect that languages such as Spanish, French, and Italian, which also use frequent function words to support grammatical cohesion, may adapt well to this methodology as they share structural similarities with English. However, challenges arise with languages that feature richer morphological complexity, like Finnish or Turkish, where grammatical meaning is embedded within word forms.
9 Conclusion
In this work we proposed an algorithm for the identification of a set of easily-inferable terms based on the context according to the desired cost-savings by using masking together with language models. We have proposed a pipeline that allows cost savings in LLM-generation by omitting the generation of terms that do not contribute to the semantic meaning, and can be easily inferred a posteriori with a smaller language model. We have evaluated our approach in the question-answering knowledge domain, where we obtained a subset of functions words that are easily inferable. We have trained 4 different language models in the task of reconstructing a distilled answer. These models obtain very good performance across all metrics without noticeable losses, while saving 20.58% tokens with respect to the original full answer. We also showed that, while the output generation is more expensive than processing input prompts, in the vast majority of cases it will be better of to process knowledge answers through the cost-saving pipeline.
10 Limitations
Although the experiments explored the capability of an LLM to generate output without including terms that could be inferred later, it did not consider how this might negatively impact the LLM’s performance. LLMs are trained on vast datasets of human-written text that follow specific grammatical rules. Consequently, their output is based on this training data and on the sequence of tokens already produced during a session. If the usual grammatical structure is absent at generation and at the previous generated data, this could disrupt the ongoing token generation.
In a proper implementation, the potential penalty would only affect the generation of individual paragraphs in a response. After generating a paragraph, the reconstruction model would restore its grammatical structure. The distilled paragraph in the context window would then be replacer by the reconstructed one. For subsequent generations, as the intrinsic grammatical structure was restored, the LLM would not face any penalties regarding the previous generations.
Moreover, in this study we have primarily focused on the knowledge domain, where structured and factual content enables effective token-saving by omitting low-information words with minimal impact on semantic clarity. However, to generalize this methodology, it is necessary to explore its applicability in other domains with more variable language patterns, such as conversational dialogue where contextual dependencies and linguistic complexity could pose challenges. Testing across these diverse domains would provide a deeper understanding of the method’s adaptability and potential adjustments needed to preserve coherence and precision in different types of language processing tasks.
Finally, another relevant limitation is that be have used automated metrics such as SacreBLEU, ROUGE, and perplexity. This provide valuable insights into grammatical similarity, semantic coherence, and fluency, but they may not fully capture the nuances of human comprehension and acceptability Warstadt (2019), especially in reconstructed texts. A future human evaluation would be desired to assess subtle aspects like readability, naturalness, and the overall quality of the reconstructed responses, which metrics alone may overlook. Including human evaluators would enable a more comprehensive understanding of the method’s effectiveness, particularly in scenarios where slight changes in phrasing or word choice could impact clarity and user satisfaction.
Disclaimer
This paper was prepared for informational purposes by the Artificial Intelligence Research group of JPMorgan Chase & Co. and its affiliates ("JP Morgan”) and is not a product of the Research Department of JP Morgan. JP Morgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.
© 2024 JPMorgan Chase & Co. All rights reserved
Appendix A Supplementary Material
A.1 Exploratory experimentation on the LLM capabilities of ommiting semantic-irrelevant terms
For this exploratory experiment we have used the model "GPT-4o" OpenAI through its API in its version "gpt-4o-2024-08-06" OpenAI (2024). The list of articles and conjunctions used are: the, a, an, as, though, because, before, even, if, that, since, so, than, unless, until, when, whenever, where, whereas, wherever, while, and, but. The set of prompts used in this experimentation are the following:
Prompt 1: Respond just in a paragraph. $\{knowledge\_question\}$
Prompt 2: Respond just in a paragraph in a distilled way removing the redundant and semantic-irrelevant terms of the language such as: $\{list\_of\_terms\}$. That is to say, omit the terms of the ideal answer that do not affect directly to the semantic meaning and can be reconstructed later on based on the context. This task will be carried out by another language model in charge of generating those omitted terms based on the list provided and reconstructing the ideal answer. Return only the distilled answer, nothing else. $\{knowledge\_question\}$
A.2 DKA Dataset elaboration
We have used the model "GPT-4o" OpenAI through its API in its version "gpt-4o-2024-08-06" OpenAI (2024). The 30 general knowledge topics, 20 concepts of each of them and 5 keywords relevant to each concept have been generated using the same LLM with a human curation after generation to verify its validity. Next we show an example of the template questions generation of one keyword (decentralization) associated to one of the concepts (blockchain) related to one of the topics (technology).
  • What is decentralization in the context of blockchain?
  • Who are the main contributors of decentralization in the context of blockchain and a summary of their contribution?
  • Define in what techniques or research has decentralization influenced and how, out of the context of blockchain
  • What is the state of the art and future or expected development of decentralization in the context of blockchain?
  • Summarize the concepts relevant for the general knowledge and understanding of decentralization in the context of blockchain?
  • What is the difference between decentralization and smart contracts in the context of blockchain?
BART T5 Qwen2-0.5 Qwen2-1.5
Training size 100.000 100.000 100.000 100.000
Epochs 5 5 3 2.5
Learning Rate 5e-5 5e-5 4e-5 3e-5
Batch Size 8 8 4 4
Gradient accumulation step 8 16 32 8
Weight decay 0.01 0.01 0.1 0.1
FP16 True True True True
Training Time (hours) 1.5 3 7.5 17
Table 4: Summary of the training process for the selected models.
Bart T5 Qwen2.5-0.5 Qwen2.5-1.5 Distilled Text
$\Theta$ Precision (%) 88.44(12.67) 88.50(12.64) 80.30(14.96) 83.60(14.07) 100.00(00.00)
$\Theta$ Recall (%) 81.58(14.74) 81.61(14.93) 82.23(14.42) 86.14(13.02) 00.00(00.00)
$\Theta$ F1 (%) 84.33(12.71) 84.39(12.85) 80.73(13.62) 84.40(12.57) 00.00(00.00)
SacreBLEU (%) 87.02(8.77) 88.38(8.53) 86.01(9.06) 88.51(8.28) 51.00(13.22)
METEOR (%) 94.87(4.57) 95.55(4.19) 95.77(3.78) 96.61(3.30) 76.10(7.11)
ROUGE-1 (%) 96.61(2.89) 96.92(2.77) 96.19(3.10) 96.79(2.85) 86.19(4.38)
ROUGE-L (%) 95.81(3.46) 96.14(3.30) 94.99(3.87) 95.94(3.46) 86.19(4.38)
Cosine Similarity (%) 99.99(00.04) 99.99(00.03) 99.98(00.03) 99.98(00.04) 99.92(00.07)
Perplexity 42.91(31.48) 39.35(29.46) 40.52(29.48) 39.01(27.76) 208.21(200.66)
Perplexity Original 37.28(26.13) 37.28(26.13) 37.28(26.13) 37.28(26.13) 37.28(26.13)
Saved Tokens (%) 18.85(5.87) 18.85(5.87) 18.85(5.87) 18.85(5.87)
Training Time (h) 1.5 3 7.5 18
Parameters (B) 0.139 0.223 0.5 1.5
Table 5: Average (standard deviation) metric results over the evaluation dataset from the Wikipedia dump.
A.3 Consistency among responses in LLMs
We have also carried out an experimentation on the consistency among different LLM answers which conditions the evaluation of this paper when dealing with generative models. We have taken a set of 20 knowledge questions, getting the answer of each of the questions 10 times in different sessions (with temperature 0 looking for maximum replicability), and compared the successive generations with the first one. As can be seen in the Table 6, although the semantic content measured by cosine similarity gives a perfect result, the other metrics related to n-grams and grammatical structure are weaker. This indicates that among the responses the semantic content is maintained, but their grammatical structure varies considerably.
Mean Variance
BLEU 0.77 0.03
ROUGE-1 0.87 0.01
ROUGE-L 0.83 0.02
METEOR 0.83 0.02
Cosine Similarity 1.00 0.00
Table 6: Exploratory results on the consistency among answers to the same questions by LLM.